40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7166     Entry time: Mon Aug 13 21:47:30 2012
Author: Yaakov 
Type: Update 
Category: STACIS 
Subject: Two changes to STACIS noise budget 

In eLog 7148 (http://nodus.ligo.caltech.edu:8080/40m/7148), Koji pointed out that the op-amp and SR560 noise values (which I took from specs and then multiplied by the geophone calibration factor to get m/s/rtHz) were waaay too low. My error was an extra multiplication factor in the plotting script.

The other change was recalculating the ADC noise by splitting a signal into two ADC channels and subtracting the time series (then taking the PSD and converting to m/s/rtHz). It compares well to the value I got by terminating the ADC channels, which was the ADC noise line in my last eLog.

Both these changes are included in the below plot:

noise_budget_8-13.bmpnoise_budget_8-13.fig

Attachment 1: noise_budget_8-13.bmp  230 kB
ELOG V3.1.3-